OpenStudy (anonymous):

If y is defined implicitly as a function of x by the equation x+(ln(y))^2 find dy/dx at the point (0,e)

OpenStudy (anonymous):

here is an example from offline, not much but should help you figure it out

OpenStudy (anonymous):

which represents a circle of radius five centered at the origin. Suppose that we wish to find the slope of the line tangent to the graph of this equation at the point (3, -4) . How could we find the derivative of y in this instance ? One way is to first write y explicitly as a function of x . Thus, x2 + y2 = 25 , y2 = 25 - x2 , and $ y = \pm \sqrt{ 25 - x^2 } $ , where the positive square root represents the top semi-circle and the negative square root represents the bottom semi-circle. Since the point (3, -4) lies on the bottom semi-circle given by $ y = - \sqrt{ 25 - x^2 } $ , the derivative of y is $ y' = - (1/2) \big( 25 - x^2 \big)^{-1/2} (-2x) = \displaystyle{ x \over \sqrt{ 25 - x^2 } } $ , i.e., $ y' = \displaystyle{ x \over \sqrt{ 25 - x^2 } } $ . Thus, the slope of the line tangent to the graph at the point (3, -4) is $ m = y' = \displaystyle{ (3) \over \sqrt{ 25 - (3)^2 } } = \displaystyle{ 3 \over 4 } $ . Unfortunately, not every equation involving x and y can be solved explicitly for y . For the sake of illustration we will find the derivative of y WITHOUT writing y explicitly as a function of x . Recall that the derivative (D) of a function of x squared, (f(x))2 , can be found using the chain rule : $ D \{ ( f(x) )^2 \} = 2 f(x) \ D \{ f(x) \} = 2 f(x) f'(x) $ . Since y symbolically represents a function of x, the derivative of y2 can be found in the same fashion : $ D \{ y^2 \} = 2 y \ D \{ y \} = 2 y y' $ . Now begin with x2 + y2 = 25 . Differentiate both sides of the equation, getting D ( x2 + y2 ) = D ( 25 ) , D ( x2 ) + D ( y2 ) = D ( 25 ) , and 2x + 2 y y' = 0 , so that 2 y y' = - 2x , and $ y' = \displaystyle{ - 2x \over 2y } = \displaystyle{ - x \over y } $ , i.e., $ y' = \displaystyle{ - x \over y } $ . Thus, the slope of the line tangent to the graph at the point (3, -4) is $ m = y' = \displaystyle{ - (3) \over (-4) } = \displaystyle{ 3 \over 4 } $ . This second method illustrates the process of implicit differentiation. It is important to note that the derivative expression for explicit differentiation involves x only, while the derivative expression for implicit differentiation may involve BOTH x AND y . The following problems range in difficulty from average to challenging

OpenStudy (anonymous):

bye did it hep?

OpenStudy (anonymous):

I got \[\frac{ -e + e ^{3xy} }{2 }\]

OpenStudy (anonymous):

do you have any idea whether I did that correctly or not?

hartnn (hartnn):

i don't see any equation in the question x+(ln(y))^2 is an expression

OpenStudy (anonymous):

sorry it should be

OpenStudy (anonymous):

\[x + (\ln y )^{2}-e ^{xy}=0\]

hartnn (hartnn):

do you know product and chain rules ?

OpenStudy (anonymous):

yeah I can show you a picture of how I did it?

hartnn (hartnn):

sure

OpenStudy (anonymous):

ok 1 sec